### v2

@@ Abstract
We study efficient mechanism design for allocating multiple heterogeneous objects. The aim is to maximize the residual surplus, the total value generated from an allocation minus the costs of screening. We discover a robust trend indicating that no-screening mechanisms, such as serial dictatorship with exogenous priority order, tend to perform better as the variety of goods increases. We analyze the underlying reasons by characterizing asymptotically efficient mechanisms in a stylized environment. We also apply an automated mechanism design approach to numerically derive efficient mechanisms and validate the trend in general environments. Building on these implications, we propose the register-invite-book system (RIB) as an efficient system for scheduling vaccination against pandemic diseases.
@@ 1 Introduction
Despite its importance, the literature has largely solved problems only in single-dimensional settings with homogeneous goods (Hartline2008). Indeed, real-world problems often involve multiple heterogeneous goods. The scheduling of vaccine appointments during the COVID-19 pandemic was a prominent example—reservation slots, each of which specifies when, where, and which vaccine will be administered, are highly heterogeneous goods, and potential recipients have diverse preferences over slots. In such settings, agents’ effort influences not only whether they secure a good but also which good they obtain. Since the efficient allocation of a large number of heterogeneous goods has not been sufficiently established in the literature, economists were unable to provide effective guidance on vaccine appointment scheduling. Consequently, various uncoordinated methods were adopted worldwide, often leading to confusion. [^2] As detailed in Section 5.2, numerous regions that adopted FCFS, including Florida, Massachusetts, Washington, D.C., and many local governments in Japan, experienced severe competition for reservation slots. By contrast, several countries and regions that adopted other systems (such as an invitation-based system) experienced far fewer such disruptions.
@@
This paper explores efficient screening mechanisms for allocating heterogeneous objects. Agents have multi-dimensional preferences over objects but only demand one item. They can signal their preferences through costly effort, which serves as a form of payment. The central planner aims to maximize residual surplus. [^3] The environment is the assignment market originally studied by koopmans1957assignment and shapley1971assignment, but our objective differs in that it is the residual surplus, where monetary transfers are not zero-sum but burned. Two prominent candidates are serial dictatorship with exogenous order (SD), where agents sequentially choose their most preferred available goods without paying screening costs, and the Vickrey-Clarke-Groves mechanism (VCG), which achieves allocative efficiency at the expense of high screening costs. Between these extremes, the planner has many options for balancing screening costs and allocative efficiency. Characterizing optimal mechanisms in this multi-dimensional setting is widely believed to be challenging.
@@
This study uncovers a robust trend: as the variety of goods increases, no-screening mechanisms like SD tend to become more efficient. Figure 1(a) illustrates this result. We consider a scenario where $2K$ agents each demand a single unit among $K$ heterogeneous objects, and each agent $i$’s value for object $k$, $v_{k}^{i}$, is drawn i.i.d. from a Weibull distribution with parameter 0.8. The horizontal axis of the figure represents the value of $K$, and the vertical axis shows the residual surplus per agent. Under this distribution, when $K=1$ (i.e., the single-good case), the optimality of VCG has been proven by Hartline2008. However, for $K\geq 3$, SD outperforms VCG, and when $K\geq 6$, the performance gap between SD and the mechanism designed by RegretNet (dutting2019optimal), an automated mechanism design method for discovering efficient mechanisms, becomes negligible. As $K$ increases further, the performance of SD continues to improve, whereas the performance gains of VCG taper off. This paper’s findings suggest that such trends will likely hold in a wide range of environments.
@@
We analyze the reduced single-dimensional environment and the distribution of the largest order statistic to characterize efficient mechanisms in the stylized environment. First, we prove that in a single-dimensional environment, a no-screening mechanism is efficient if and only if the value distribution satisfies the new better than used in expectation (NBUE) property. NBUE is a weaker condition than increasing hazard rate (IHR), shown as a sufficient condition by Hartline2008. Second, we prove that, for any marginal value distribution, as the variety of goods, $K$, increases, the distribution of the largest order statistic tends to satisfy NBUE and IHR, implying that no-screening mechanisms become more likely to be efficient as the number of goods grows. Third, using extreme value theory, we analyze the limit as the variety of goods $K$ approaches infinity. We show that in this limit, the no-screening mechanism is efficient for a broad class of marginal distributions, and even when it is not, it is practically close to being efficient.
@@
Figure 1(b) illustrates the observation for the continuous market. Parallel to the setting of Figure 1(a), there are $K$ types of objects with an equal mass and agents whose mass is twice that of the objects. Each agent $i$’s value for each object type $k$, $v_{k}^{i}$, is independently drawn from a Weibull distribution with parameter 0.6. We observe that the performance of SD improves as $K$ increases. For such a continuous market, we can also prove as a theorem that whenever the marginal value distribution satisfies the von Mises condition with parameter 0 (VonMises1936), such as the Weibull distribution (with any parameter), the performance ratio of SD to the exact efficient mechanism converges to one. (The setting and results will be reviewed in Section 3.9.)
@@
Analytical and even numerical analyses for finite markets with heterogeneous objects have been known to be extremely challenging. Nevertheless, we leverage deep-learning techniques for automated mechanism design, which have recently been developed with a primary focus on revenue maximization (dutting2019optimal), to derive mechanisms for maximizing residual surplus. Consistent with the results from the analysis of stylized environments, we show that in these finite markets, as the number of objects (or types of objects) increases, the performance of SD improves, as depicted in Figure 1(a).
@@
We further conduct numerical analyses of cases with correlated values, which are practically important but challenging for the application of extreme value theory. [^4] According to a review by Majumdar2020, except for cases with weakly correlated values—where the correlation length is finite and a renormalization group-like argument can reduce the problem to that of uncorrelated values—the behavior of the extreme value statistic in the limit is largely unknown. Correlation between values assigned to different objects by a single agent (within-agent correlation) brings the situation closer to a single-dimensional (i.e., homogeneous-good) environment, thereby weakening the effect that the multiplicity of objects improves the performance of no-screening mechanisms. However, the correlation between values assigned to a single object by different agents (between-agent correlation) enhances the relative performance of no-screening mechanisms, as in such cases, the screening costs outweigh the allocative benefits. While these types of correlations can either amplify or mitigate the effects of increasing the variety of goods, they do not qualitatively impact the policy implication that the relative performance of SD improves as the variety of goods increases.
@@ 2 Related Literature
Maximizing residual surplus parallels designing mechanisms with monetary transfers, such as auctions that maximize welfare while discarding payments. This problem, often called money burning, has been studied extensively. For environments with homogeneous goods, (Hartline2008; yoon2011optimal; condorelli2012money; chakravarty2013optimal) provide characterizations of efficient mechanisms. More recent works, such as Akbarpour2023comparison and tokarski2025equitable, combine wasteful effort (ordeals) with other screening devices, such as payments or damages, to propose mechanisms that screen not only for the need for the good but also simultaneously for other private characteristics, including social welfare weights and the cost of completing the ordeal.
@@
However, characterizing optimal mechanisms for multi-dimensional types with multiple objects remains challenging. While revenue maximization has received considerable attention (e.g., manelli2006bundling; pavlov2011optimal; Haghpanah2020bundling; giannakopoulos2018duality; daskalakis2015strong), no complete characterization exists, even in simple scenarios. Similarly, maximizing residual surplus in multi-object environments is complex. The existing literature provides only partial characterizations, and typically only in restricted environments. For example, tokarski2025damages shows that when distributing two goods using deterministic mechanisms, one can improve residual surplus by deliberately degrading the value of the goods. Fotakis2016EfficientMB and goldner2024simple propose approximation mechanisms with performance guarantees. Although our characterization is also confined to a stylized environment, our results suggest that when the number and variety of goods are large, the efficient mechanism may become relatively simple and practically implementable.
@@
Reflecting the analytical difficulty of designing mechanisms for environments with multi-dimensional types, methods for automated mechanism design have also advanced substantially (sandholm2003automated; conitzer2004self; dutting2019optimal). To our knowledge, our study is the first to apply these methods to residual surplus maximization. It is also worth noting that our study provides practical insights, whereas revenue-maximizing mechanisms are often impractical for real-world implementation because they frequently have complex structures or are highly sensitive to the environment.
@@
Serial dictatorship (SD) mechanisms are well-studied for their desirable properties. Pycia2024random demonstrate that the random serial dictatorship in sequential form is the only mechanism that satisfies Pareto efficiency, symmetry, and obvious strategy-proofness (li2017obviously). bade2015serial demonstrate SD’s robustness when agents can acquire preference information, and hakimov2023costly confirm its empirical effectiveness in school choice. Additionally, SD achieves at least 11/e1-1/e of the first-best matching size (Krysta2014size; Noda2020size) and prevents scalping (Hakinov2021appointment). Combining these advantages with our findings, SD is a promising option for allocating many heterogeneous goods without payments.
@@
Finally, research on medical resource allocation during crises has grown rapidly. Studies have analyzed incentives for production (Ahuja2021preparing; castillo2021market; Athey2022expanding), capacity design (Noda2018large), priorities (Gans2022allocation; Akbarpour2023prioritization), and reserves (Pathak2022reserve; Pathak2023Vaccine). Our paper complements this literature by focusing on the downstream process of vaccine distribution, offering practical insights for future public health crises.
@@ 3.1 Model
Throughout this section, we consider direct mechanisms that determine an allocation and payment based only on the agent’s own report and the distribution of the reported type profiles. [^5] This property is obtained in the continuous limit of a sequence of mechanisms satisfying liu2016ordinal’s (liu2016ordinal) regularity condition. This restriction has two effects. First, we can drop an index representing the agent’s identity because an agent’s identity does not alter the agent’s outcome. Second, in a continuous market, there is no aggregate uncertainty, and thus the distribution of reported types remains constant in equilibrium. Accordingly, we can also omit the other agents’ type reports (or their distribution) from the arguments of the mechanism. A mechanism $\mathcal{M}$ is comprised of an allocation rule $\mathbf{x}:V\to[0,1]^{K}$ and a payment rule $p:V\to\mathbb{R}_{+}$, where $\mathbf{x}(\mathbf{v})$ and $p(\mathbf{v})$ are the respective allocation and payment when the agent has a valuation $\mathbf{v}$. Let $F:V\to[0,1]$ be the cumulative distribution function of agents’ valuations. Our main measure of social welfare, residual surplus from a mechanism $\mathcal{M}=(\mathbf{x},p)$ is defined as
@@ 3.2 Additional Restrictions for Analytical Tractability
The proof is similar to that of Theorem 1 of rahme2021permutation.
@@ 3.3 Reduction to a Single-Dimensional Environment
Proofs are shown in Appendix B. This theorem holds only in the continuous i.i.d. market. When the popularity or scarcity of objects is asymmetric, some agents may, of course, end up receiving unpopular or abundant objects, even if those are not their favorite ones. In finite markets, objects symmetric ex ante may become asymmetric ex post, depending on realized preferences, causing some agents to receive their second or lower choices. However, in broad environments, agents tend to receive highly preferred objects when both goods and preferences are diverse. Theorem 1 captures this insight in its most extreme yet tractable form. As we later show through numerical analysis, the qualitative implications derived from the analysis of the continuous i.i.d. market carry over to more general environments.
@@ 3.4 Mechanisms
The revenue equivalence theorem of myerson1981optimal implies that the gross surplus $\bar{m}\mathbb{E}[v]$ cannot be improved further without burning payments.
This result is achieved in a continuous i.i.d. market if we apply serial dictatorship (SD), in which agents take their most preferred objects sequentially following an exogenous priority order. In SD, the first $\bar{m}$ agents of the priority order, facing all goods available, obtain their favorite goods. Exactly when mass $\bar{m}$ agents finish choosing, all goods are simultaneously exhausted, and the remaining $1-\bar{m}$ agents receive nothing. [^8] Formally, the outcome of SD in a continuous environment is specified by an eating mechanism. While we skip it because this paper does not intensively analyze the general continuous multidimensional environment, interested readers should refer to Noda2018large. The asymptotic equivalence of (random) SD and an eating mechanism is established by che2010asymptotic.
@@
The revenue equivalence theorem of myerson1981optimal implies that to achieve the gross surplus $\bar{m}\mathbb{E}[v|v\geq q]$, each allocated agent needs to burn $q$; thus, the total expected burned payment is $\bar{m}q$.
For general environments, including finite environments and environments with continuous agents and objects but non-i.i.d. values, the Vickrey-Clarke-Groves mechanism (VCG), which requires each agent to pay their externality, achieves allocative efficiency. Furthermore, green1977characterization; green1979incentives; holmstrom1979groves show that, under certain conditions, VCG is the unique direct mechanism that achieves allocative efficiency even in environments with multi-dimensional types.
Direct mechanisms such as VCG are not practically used in money-burning problems. However, when indirect mechanisms are employed, the resulting equilibrium allocations often coincide with or resemble those achieved by direct mechanisms. For example, during the COVID-19 pandemic, FCFS was used to allocate vaccine reservation slots. Under this mechanism, agents who claim after the distribution begins receive, in order of arrival, their most preferred goods among those still available. If we formulate this as a problem where agents freely choose their payment (effort cost) to claim earlier, then (i) agents who pay a higher payment are more likely to obtain goods, and (ii) payments are made regardless of whether the good is obtained. In a reduced single-dimensional market, this incentive structure is equivalent to an all-pay auction. In the reduced environment, an all-pay auction implements the same allocation rule and expected utilities as the VCG by revenue equivalence. Furthermore, whereas FCFS might yield a different outcome from VCG in general environments, VCG, which achieves allocative efficiency, represents the opposite extreme of a no-screening mechanism and serves as a useful benchmark. [^9] In the finite market studied in Section 4, we assume finitely many agents and objects, unit demand, and quasi-linear utility. This setting coincides with the assignment markets by koopmans1957assignment; shapley1971assignment. In this environment, it is also known from crawford1981job; demange1986multi that dynamic auctions, which gradually raise the prices of over-demanded goods, can find the allocation that maximizes gross surplus.

### v3

@@ Abstract
We study efficient mechanism design for allocating multiple heterogeneous objects. The aim is to maximize the residual surplus, the total value generated from an allocation minus the costs of screening. We discover a robust trend indicating that no-screening mechanisms, such as serial dictatorship with exogenous priority order, tend to perform better as the variety of goods increases. We analyze the underlying reasons by characterizing asymptotically efficient mechanisms in a stylized environment. We also apply an automated mechanism design approach to numerically derive efficient mechanisms and validate the trend in general environments. Building on these implications, we propose the register-invite-book system (RIB) as an efficient system for scheduling vaccinations against pandemic diseases.
@@ 1 Introduction
Despite its importance, the literature has largely solved problems only in single-dimensional settings with homogeneous goods (Hartline and Roughgarden, 2008). Indeed, real-world problems often involve multiple heterogeneous goods. The scheduling of vaccine appointments during the COVID-19 pandemic was a prominent example—reservation slots, each of which specifies when, where, and which vaccine will be administered, are highly heterogeneous goods, and potential recipients have diverse preferences over slots. In such settings, agents’ effort influences not only whether they secure a good but also which good they obtain. Since the efficient allocation of a large number of heterogeneous goods has not been sufficiently established in the literature, economists were unable to provide effective guidance on vaccine appointment scheduling. Consequently, various uncoordinated methods were adopted worldwide, often leading to confusion. [^2] As detailed in Section 5.2, numerous regions that adopted FCFS, including Florida, Massachusetts, Washington, D.C., and many local governments in Japan, experienced severe competition for reservation slots. By contrast, several countries and regions that adopted other systems (such as an invitation-based system) experienced far fewer such disruptions.
@@
This paper explores efficient screening mechanisms for allocating heterogeneous objects. Agents have multi-dimensional preferences over objects but only demand one item. They can signal their preferences through costly effort, which serves as a form of payment. The central planner aims to maximize residual surplus. [^3] The environment is the assignment market originally studied by Koopmans and Beckmann (1957) and Shapley and Shubik (1971), but our objective differs in that it is the residual surplus, where monetary transfers are not zero-sum but burned. Two prominent candidates are serial dictatorship with exogenous order (SD), where agents sequentially choose their most preferred available goods without paying screening costs, and the Vickrey-Clarke-Groves mechanism (VCG), which achieves allocative efficiency at the expense of high screening costs. Between these extremes, the planner has many options for balancing screening costs and allocative efficiency. Characterizing optimal mechanisms in this multi-dimensional setting is widely believed to be challenging.
@@
This study uncovers a robust trend: as the variety of goods increases, no-screening mechanisms like SD tend to become more efficient. Figure 1(a) illustrates this result. We consider a scenario where $2K$ agents each demand a single unit among $K$ heterogeneous objects, and each agent $i$’s value for object $k$, $v_{k}^{i}$, is drawn i.i.d. from a Weibull distribution with parameter 0.8. The horizontal axis of the figure represents the value of $K$, and the vertical axis shows the residual surplus per agent. Under this distribution, when $K=1$ (i.e., the single-good case), the optimality of VCG has been proven by Hartline and Roughgarden (2008). However, for $K\geq 3$, SD outperforms VCG, and when $K\geq 6$, the performance gap between SD and the mechanism designed by RegretNet (Dütting et al., 2019), an automated mechanism design method for discovering efficient mechanisms, becomes negligible. As $K$ increases further, the performance of SD continues to improve, whereas the performance gains of VCG taper off. This paper’s findings suggest that such trends will likely hold in a wide range of environments.
@@
We analyze the reduced single-dimensional environment and the distribution of the largest order statistic to characterize efficient mechanisms in the stylized environment. First, we prove that in a single-dimensional environment, a no-screening mechanism is efficient if and only if the value distribution satisfies the new better than used in expectation (NBUE) property. NBUE is a weaker condition than increasing hazard rate (IHR), shown as a sufficient condition by Hartline and Roughgarden (2008). Second, we prove that, for any marginal value distribution, as the variety of goods, $K$, increases, the distribution of the largest order statistic tends to satisfy NBUE and IHR, implying that no-screening mechanisms become more likely to be efficient as the number of goods grows. Third, using extreme value theory, we analyze the limit as the variety of goods $K$ approaches infinity. We show that in this limit, the no-screening mechanism is efficient for a broad class of marginal distributions, and even when it is not, it is practically close to being efficient.
@@
Figure 1(b) illustrates the observation for the continuous market. Parallel to the setting of Figure 1(a), there are $K$ types of objects with an equal mass and agents whose mass is twice that of the objects. Each agent $i$’s value for each object type $k$, $v_{k}^{i}$, is independently drawn from a Weibull distribution with parameter 0.6. We observe that the performance of SD improves as $K$ increases. For such a continuous market, we can also prove as a theorem that whenever the marginal value distribution satisfies the von Mises condition with parameter 0 (von Mises, 1936), such as the Weibull distribution (with any parameter), the performance ratio of SD to the exact efficient mechanism converges to one. (The setting and results will be reviewed in Section 3.9.)
@@
Analytical and even numerical analyses for finite markets with heterogeneous objects have been known to be extremely challenging. Nevertheless, we leverage deep-learning techniques for automated mechanism design, which have recently been developed with a primary focus on revenue maximization (Dütting et al., 2019), to derive mechanisms for maximizing residual surplus. Consistent with the results from the analysis of stylized environments, we show that in these finite markets, as the number of objects (or types of objects) increases, the performance of SD improves, as depicted in Figure 1(a).
@@
We further conduct numerical analyses of cases with correlated values, which are practically important but challenging for the application of extreme value theory. [^4] According to a review by Majumdar et al. (2020), except for cases with weakly correlated values—where the correlation length is finite and a renormalization group-like argument can reduce the problem to that of uncorrelated values—the behavior of the extreme value statistic in the limit is largely unknown. Correlation between values assigned to different objects by a single agent (within-agent correlation) brings the situation closer to a single-dimensional (i.e., homogeneous-good) environment, thereby weakening the effect that the multiplicity of objects improves the performance of no-screening mechanisms. However, the correlation between values assigned to a single object by different agents (between-agent correlation) enhances the relative performance of no-screening mechanisms, as in such cases, the screening costs outweigh the allocative benefits. While these types of correlations can either amplify or mitigate the effects of increasing the variety of goods, they do not qualitatively impact the policy implication that the relative performance of SD improves as the variety of goods increases.
@@ 2 Related Literature
Maximizing residual surplus parallels designing mechanisms with monetary transfers, such as auctions that maximize welfare while discarding payments. This problem, often called money burning, has been studied extensively. For environments with homogeneous goods, (Hartline and Roughgarden, 2008; Yoon, 2011; Condorelli, 2012; Chakravarty and Kaplan, 2013) provide characterizations of efficient mechanisms. More recent works, such as Akbarpour et al. (2023b) and Tokarski (2025a), combine wasteful effort (ordeals) with other screening devices, such as payments or damages, to propose mechanisms that screen not only for the need for the good but also simultaneously for other private characteristics, including social welfare weights and the cost of completing the ordeal. Dworczak (2025) studies the allocation of money under costly screening and derives conditions under which ordeal mechanisms outperform lump-sum transfers due to redistributive concerns. Yang (2024) shows that when the productive and cost components of agents’ types are positively correlated, an optimal mechanism does not use ordeals.
@@
However, characterizing optimal mechanisms for multi-dimensional types with multiple objects remains challenging. While revenue maximization has received considerable attention, complete characterization is analytically intractable, even in simple scenarios (Thanassoulis, 2004; Pycia, 2006; Manelli and Vincent, 2006). Because full characterizations in general environments are elusive, the literature has focused on tractable special environments (McAfee and McMillan, 1988; Pavlov, 2011; Daskalakis et al., 2015, 2017; Giannakopoulos and Koutsoupias, 2018; Haghpanah and Hartline, 2021; Bikhchandani and Mishra, 2022; Ghili, 2023; Yang, 2025), approximately optimal mechanisms (Cai et al., 2012, 2016, 2019; Hart and Nisan, 2017), and worst-case optimal mechanisms (Carroll, 2017; Deb and Roesler, 2021, 2023; Che and Zhong, 2021, 2024). Parallel to revenue maximization, maximizing residual surplus in multi-object environments is complex. The existing literature provides only partial characterizations, and typically only in restricted environments. For example, Tokarski (2025b) shows that when distributing two goods using deterministic mechanisms, one can improve residual surplus by deliberately degrading the value of the goods. Fotakis et al. (2016) and Goldner and Lundy (2024) propose approximation mechanisms with performance guarantees. Although our characterization is also confined to a stylized environment, our results suggest that when the number and variety of goods are large, the efficient mechanism may become relatively simple and practically implementable.
@@
Reflecting the analytical difficulty of designing mechanisms for environments with multi-dimensional types, methods for automated mechanism design have also advanced substantially (Sandholm, 2003; Conitzer and Sandholm, 2004; Dütting et al., 2019). To our knowledge, our study is the first to apply these methods to residual surplus maximization. It is also worth noting that our study provides practical insights, whereas revenue-maximizing mechanisms are often impractical for real-world implementation because they frequently have complex structures or are highly sensitive to the environment.
@@
Serial dictatorship (SD) mechanisms are well-studied for their desirable properties. Pycia and Troyan (2024) demonstrate that the random serial dictatorship in sequential form is the only mechanism that satisfies Pareto efficiency, symmetry, and obvious strategy-proofness (Li, 2017). Bade (2015) demonstrate SD’s robustness when agents can acquire preference information, and Hakimov et al. (2023) confirm its empirical effectiveness in school choice. Additionally, SD achieves at least 11/e1-1/e of the first-best matching size (Krysta et al., 2014; Noda, 2020) and prevents scalping (Hakimov et al., 2021). Combining these advantages with our findings, SD is a promising option for allocating many heterogeneous goods without payments.
@@
Finally, research on medical resource allocation during crises has grown rapidly. Studies have analyzed incentives for production (Ahuja et al., 2021; Castillo et al., 2021; Athey et al., 2022), capacity design (Noda, 2018), priorities (Gans, 2022; Akbarpour et al., 2023a), and reserves (Pathak et al., 2022, 2024). Our paper complements this literature by focusing on the downstream process of vaccine distribution, offering practical insights for future public health crises.
@@ 3.1 Model
Throughout this section, we consider direct mechanisms that determine an allocation and payment based only on the agent’s own report and the distribution of the reported type profiles. [^5] This property is obtained in the continuous limit of a sequence of mechanisms satisfying Liu and Pycia’s (2016) regularity condition. This restriction has two effects. First, we can drop an index representing the agent’s identity because an agent’s identity does not alter the agent’s outcome. Second, in a continuous market, there is no aggregate uncertainty, and thus the distribution of reported types remains constant in equilibrium. Accordingly, we can also omit the other agents’ type reports (or their distribution) from the arguments of the mechanism. A mechanism $\mathcal{M}$ is comprised of an allocation rule $\mathbf{x}:V\to[0,1]^{K}$ and a payment rule $p:V\to\mathbb{R}_{+}$, where $\mathbf{x}(\mathbf{v})$ and $p(\mathbf{v})$ are the respective allocation and payment when the agent has a valuation $\mathbf{v}$. Let $F:V\to[0,1]$ be the cumulative distribution function of agents’ valuations. Our main measure of social welfare, residual surplus from a mechanism $\mathcal{M}=(\mathbf{x},p)$ is defined as
@@ 3.2 Additional Restrictions for Analytical Tractability
The proof is similar to that of Theorem 1 of Rahme et al. (2021).
@@ 3.3 Reduction to a Single-Dimensional Environment
Proofs are provided in Appendix B. This theorem holds only in the continuous i.i.d. market. When the popularity or scarcity of objects is asymmetric, some agents may, of course, end up receiving unpopular or abundant objects, even if those are not their favorite ones. In finite markets, objects symmetric ex ante may become asymmetric ex post, depending on realized preferences, causing some agents to receive their second or lower choices. However, in broad environments, agents tend to receive highly preferred objects when both goods and preferences are diverse. Theorem 1 captures this insight in its most extreme yet tractable form. As we later show through numerical analysis, the qualitative implications derived from the analysis of the continuous i.i.d. market carry over to more general environments.
@@ 3.4 Mechanisms
The revenue equivalence theorem of Myerson (1981) implies that the gross surplus $\bar{m}\mathbb{E}[v]$ cannot be improved further without burning payments.
This result is achieved in a continuous i.i.d. market if we apply serial dictatorship (SD), in which agents take their most preferred objects sequentially following an exogenous priority order. In SD, the first $\bar{m}$ agents of the priority order, facing all goods available, obtain their favorite goods. Exactly when mass $\bar{m}$ agents finish choosing, all goods are simultaneously exhausted, and the remaining $1-\bar{m}$ agents receive nothing. [^8] Formally, the outcome of SD in a continuous environment is specified by an eating mechanism. While we skip it because this paper does not intensively analyze the general continuous multidimensional environment, interested readers should refer to Noda (2018). The asymptotic equivalence of (random) SD and an eating mechanism is established by Che and Kojima (2010).
@@
The revenue equivalence theorem of Myerson (1981) implies that to achieve the gross surplus $\bar{m}\mathbb{E}[v|v\geq q]$, each allocated agent needs to burn $q$; thus, the total expected burned payment is $\bar{m}q$.
For general environments, including finite environments and environments with continuous agents and objects but non-i.i.d. values, the Vickrey-Clarke-Groves mechanism (VCG), which requires each agent to pay their externality, achieves allocative efficiency. Furthermore, Green and Laffont (1977, 1979); Holmström (1979) show that, under certain conditions, VCG is the unique direct mechanism that achieves allocative efficiency even in environments with multi-dimensional types.
Direct mechanisms such as VCG are not practically used in money-burning problems. However, when indirect mechanisms are employed, the resulting equilibrium allocations often coincide with or resemble those achieved by direct mechanisms. For example, during the COVID-19 pandemic, FCFS was used to allocate vaccine reservation slots. Under this mechanism, agents who claim after the distribution begins receive, in order of arrival, their most preferred goods among those still available. If we formulate this as a problem where agents freely choose their payment (effort cost) to claim earlier, then (i) agents who pay a higher payment are more likely to obtain goods, and (ii) payments are made regardless of whether the good is obtained. In a reduced single-dimensional market, this incentive structure is equivalent to an all-pay auction. In the reduced environment, an all-pay auction implements the same allocation rule and expected utilities as the VCG by revenue equivalence. Furthermore, whereas FCFS might yield a different outcome from VCG in general environments, VCG, which achieves allocative efficiency, represents the opposite extreme of a no-screening mechanism and serves as a useful benchmark. [^9] In the finite market studied in Section 4, we assume finitely many agents and objects, unit demand, and quasi-linear utility. This setting coincides with the assignment markets by Koopmans and Beckmann (1957); Shapley and Shubik (1971). In this environment, it is also known from Crawford and Knoer (1981); Demange et al. (1986) that dynamic auctions, which gradually raise the prices of over-demanded goods, can find the allocation that maximizes gross surplus.